what is log2(x-2) +log2 (x+2) = 5? Thank you.
Answer by user_dude2008(1825)

all logs are base 2

log2 (x-2) +log2 (x+2) = 5


log2[ (x-2)(x+2) ] = 5


(x-2)(x+2) = 2^5


(x-2)(x+2) = 32


x^2-4 = 32


x^2 = 36

x = -6 (extraneous) or x = 6


Ans: only solution is x = 6